# Fight like a Woman:

### v2

@@ Abstract
We investigate the impact of judges’ gender on the outcome of domestic violence cases. Using data from São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. To show that this decision gap rises due to different gender perspectives about domestic violence instead of rising due to female judges being tougher than male judges, we compare it against gender conviction rate gaps in similar types of crimes. We find that the gender conviction rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p. larger) and for other physical assault cases (8 p.p. larger). Lastly, we find evidence that at least two channels explain this gender conviction rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria.
@@ Introduction
Violence against women is frequently perpetrated inside the home by the ones closest to them. The United Nations Office on Drugs and Crime (2023) estimates that 55% of all homicides against women worldwide in 2022 were committed by intimate partners or other family members. As this violence is a criminal offense, key components of the fight against domestic violence are the Criminal Court System and how court judges approach this type of crime (Beecher-Monas, 2001). For instance, the Human Rights Watch (1991) argues that, in the 1980s and early 1990s, Brazilian judges failed to analyze crimes of domestic violence against women in a non-discriminatory manner. [^1] In the years since, Brazilian authorities have made important legal and institutional changes to address domestic violence crimes. The most prominent of these is the Maria da Penha Law, enacted in 2006, which characterizes domestic violence as a specific type of criminal offense. Moreover, Kafka et al. (2019) finds that American judges in North Carolina use gender stereotypes to argue that some women want to “game the system” with “frivolous cases” when requesting protective orders against their abusive partners.
@@
Since discriminatory gender stereotypes may vary depending on the decision maker’s individual characteristics, we investigate the impact of judges’ subjectivity on the outcome of domestic violence cases. In particular, we focus on the effect of a judge’s gender on their decisions, documenting a gender conviction-rate gap specific to domestic violence cases. We also analyze possible forces driving this difference.
@@
To do this, we collect data from São Paulo, Brazil, between 2010 and 2019 and compare conviction rates by the deciding judge’s gender after controlling for the judge’s career characteristics. Case characteristics are independent of the presiding judge’s gender because cases are randomly allocated to judges within each court district in Brazil. Hence, the gender conviction-rate gap captures different decision criteria by judges of different genders.
@@
In this analysis, we find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to gender decision gaps for two other types of crime: misdemeanors and physical assault. Because we focus on domestic violence crimes that do not result in death or long-term physical impairment, misdemeanors are comparable to domestic violence cases in their sentence severity (according to Brazilian Law). Physical assault cases are comparable to domestic violence cases because both involve some sort of physical aggression.
@@
We find that the gender conviction-rate gap for domestic violence cases is more than twice the size of that of other misdemeanors. Furthermore, while there is a gender conviction-rate gap for domestic violence cases, there are no differences between female and male judges when analyzing other types of physical assault cases. Consequently, the gender conviction-rate gap for domestic violence cases likely arises due to different perspectives about this type of crime rather than because female judges are, in general, more punitive than their male counterparts.
@@
To investigate if the salience of gender identity matters (representational account), we test if the gender difference in strictness is more pronounced in intimate partner violence cases. To do so, we explore each case’s full sentence and estimate whether female judges are more likely to convict a defendant for a domestic violence offense and simultaneously flag the case as associated with intimate partner violence. We find that the gender conviction rate gap is 37.5% higher (9 p.p.) in cases in which a female judge may identify more strongly with the victim than a male judge, while there is no gender difference in cases that are not flagged as intimate partner violence offenses. This difference is associated with the representational account of the in-group bias because the incidence and tolerance of this type of offense are strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
@@
To test for gender differences in the interpretation of the evidence provided in domestic violence cases (informational account), we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence. We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while they differ when making decisions about cases with more subjective evidence. Hence, female judges process the information contained in domestic violence cases differently from their male peers, as suggested by the informational account of the in-group bias.
Finally, we document that the difference in conviction rates between male and female judges is more pronounced later in their careers, which is when career pressures are weaker. We define late-career judges as those in the last stage of the career of a Trial Court judge. Because judges in Brazil are not elected and have job security, [^2] By law, they acquire job stability after 2 years on the job. The goal is to insulate them from political pressure. promotion is the main form of career incentive. However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages (Laneuville, 2024). We find that there is no gender conviction-rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage. This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction-rate gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
@@
This article contributes to the literature on judicial biases. Harris and Sen (2019) reviews this literature, focusing on ideology, race and gender gaps. Within the studies focusing on in-group bias, many authors find evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Boyd, Epstein, and Martin, 2010; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Similarly to our work, Kruttschnitt and Savolainen (2009), Boyd, Epstein, and Martin (2010) Bielen and Grajzl (2020), and Chen et al. (2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking) in Finland, federal appellate cases in the U.S., civil cases in Belgium, and for female offenders in Kenya, respectively. Our work differs from theirs in that we focus on a single common crime type (domestic violence) whose victims are mostly women. Furthermore, while Boyd, Epstein, and Martin (2010) finds that in-group bias in federal appellate cases in the U.S. is consistent with the informational account only, we find that in-group bias in criminal cases in Brazil is consistent with the representational and informational accounts.
@@
The paper is organized as follows. Section 2 describes our dataset, while Section 3 explains our empirical strategy. Moreover, Section 4 documents the gender conviction rate gap, while Section 5 discusses its driving forces. Finally, Section 6 argues that the difference in perspective of early-career judges may not be entirely reflected in the gender conviction-rate gap, and Section 7 concludes.
@@
We collect data from all criminal cases brought to the Justice Court System in the State of São Paulo, Brazil, between January 4th, 2010, and December 3rd, 2019. This dataset includes the date each case was assigned to a courtroom, the court district where the crime took place, the judge assigned to the case, the full text of the sentence, a list of the main events of the case, and the primary and secondary crime types associated to the case.
@@
We use the full-sentence text to construct our main outcomes. To do this, we adapt the classification algorithm developed by Possebom (2023) so that it classifies domestic violence case outcomes in a more precise manner. [^4] See Appendix A for an overview of our data-construction process. We classify sentences into one of six possible judicial decision categories.
The first decision type is conviction. This variable equals one when the judge convicts at least one of the defendants listed in the case of at least one of the charges on the case. As a consequence of conviction, the defendant receives some type of punishment (e.g., incarceration, fine, or community service) and a criminal record.
The second decision category is a non-prosecution agreement (NPA). This variable equals one when the defendant and the district attorney’s office strike a deal to exchange a criminal charge for a lighter punishment, such as a fine. Such agreements must be authorized by a judge and do not generate a criminal record.
The third decision type is dismissal. A judge may dismiss a case if they consider it a less serious offense and the defendant fulfills a series of requirements. For example, the defendant may be forced to show up in court periodically for two years and be forbidden from going to bars. Similarly to NPAs, dismissed cases do not generate criminal records.
The fourth decision category is acquittal. This variable equals one if the defendant is ruled “not guilty” on the charges listed.
Lastly, we account for cases that do not come to a conclusion because either the defendant died before sentencing or because the criminal charges expired. In the latter situation, defendants can no longer be convicted according to the statute of limitations because too much time has elapsed since the day the crime was committed.
@@
The district attorney’s office defines each case’s primary and secondary crime prior to judge assignment. We construct three samples based on those crime types.
@@
We define a ‘physical assault’ sample of cases whose primary type is physical assault. These cases are associated with other forms of assault such as street fights, bar fights, or football hooliganism. We call ‘other physical assault’ the physical assault cases that do not involve domestic violence.
Third, we define a ‘misdemeanors’ sample comprised of those crimes that carry a maximum incarceration sentence of less than four years. This threshold is relevant because, according to Brazilian Law, being convicted of these crime types leads to a criminal record, but not to incarceration. Sentences must be converted into fines or community service. As discussed by Possebom (2023) and Acerenza, Possebom, and Sant’Anna (2024), most misdemeanor crimes in Brazil are theft offenses. We call ‘other misdemeanors’ the misdemeanor offenses unrelated to domestic violence.
@@
The productivity data from the São Paulo Court of Justice website contains records on the total amount of work done by each judge in any given month and court district from January 2011 to December 2019.666For example, it measures the number of hearings or disposed cases per month for each judge in each court district. We use this information to keep track of which judges are active in each court district-month pair. [^7] Months are the natural unit of time for keeping track of activity and work volume because salaries in Brazil are paid monthly. Furthermore, most hiring, reallocation, and termination takes effect at the beginning of each month.
@@
Our final dataset is at the criminal-case level. It is crucial for our identification strategy that, for each case, the gender of the judge is randomly assigned and correctly identified.
@@
Third, since we compare decisions made by female and male judges, we require the judge’s gender to be randomly allocated to cases. In other words, the ex-ante probability of a female judge being assigned to each case must be different from zero or one, which would not be the case in courts without judges of both genders. To enforce this restriction, we limit the sample to cases from court district-month pairs with at least one male and one female active judge in the productivity dataset.
Last, we need to ensure that the assigned judge’s gender is determined correctly. To do this, we focus on cases allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment. [^9] Cases are randomly allocated to courtrooms and instead of directly to individual judges. If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
Using this final sample, we compute summary statistics for each possible case outcome and crime type. Table 1 displays case outcomes in the columns and crime types in the rows. It also reports the share of female judges analyzing each type of crime. Domestic violence cases are more frequently acquitted than other misdemeanors or physical assault cases. As conviction rates are similar across types of crimes, this higher acquittal is due to fewer NPAs and dismissed cases. We also note that the share of female judges is similar across crime types and that there are 270 female judges and 425 male judges in our entire sample.
Table 1: Case Outcomes and Share of Female Judges by Crime Type in %
Crime Conviction NPA Acquittal Dismissal Expired Dead Female Obs
@@
DV 32.5 1.1 51.4 4.3 8.8 2.0 32.1 4,358
Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 36.4 39,815
Other PA 27.2 7.0 34.1 20.6 10.4 1.9 38.1 1,648
  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Mis), and physical assault cases that are not domestic violence (PA). The first six columns indicate the frequency of each type of case outcome. The conviction dummy equals one when at least one defendant is convicted in at least one charge. The NPA dummy equals one when the defendant and the district attorney’s office sign an NPA. The dismissal dummy equals one if the case is dismissed. The acquittal dummy equals one if the defendant is found not guilty. The “expired” dummy equals one if at least one of the charges in the case has expired. The “Dead” dummy equals one if the defendant died before the sentence’s date. The rows do not sum to 100% because cases with multiple charges can have different outcomes for each charge. The second to last column shows the share of female judges in each sample while the last column shows the size of each sample.
@@
We measure the gender conviction rate gap for domestic violence cases by comparing the conviction probability for similar cases allocated to female and male judges. We do this by running a linear regression model on the sample of domestic violence crimes:
@@
where $\text{Convicted}_{it}$ indicates whether the cases resulted in a conviction, $\text{Female Judge}_{it}$ indicates whether the case was assigned to a female judge, $X_{it}$ are covariates including case characteristics (an indicator for whether the defendant was already in jail and an indicator for whether the police caught them committing the crime) and pre-determined judge characteristics (career stage, total experience, and experience in their current career stage as of the day the case was assigned), and $\text{Court District}_{i}\times\text{Quarter}_{t}$ is a full set of court district-quarter pair fixed effects.
@@
Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence. To measure whether this conviction-rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types. Now, our sample consists of domestic violence cases and one type of comparable criminal case. We estimate the following linear regression model:
@@
where $\text{Domestic Violence}_{it}$ indicates whether the case is associated with domestic violence. We use two categories of comparable crime types: (i) offenses with similar sentences (misdemeanors) and (ii) offenses of a similar nature (physical assault).
In Equation (2), our parameter of interest is $\beta_{3}$. This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of crimes of comparable severity (coefficient $\beta_{2}$) and for the possibility that female judges are stricter than male judges in general (coefficient $\beta_{1}$). Consequently, this coefficient captures the part of the gender conviction-rate gap that is likely due to different gender perspectives about domestic violence cases specifically. For this reason, we interpret the $\beta_{3}$ coefficient as capturing a type of in-group bias according to the definition given by Shayo and Zussman (2011). We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) may act differently from members of the other group (male judges) when analyzing this type of crime in particular.
This interpretation for the $\beta_{3}$ coefficient is valid if the judge’s gender is independent of the case’s potential outcome after conditioning on the type of crime. This exogeneity assumption is again valid according to Brazilian Law since cases are randomly allocated to judges within court districts and crime types are determined by the district attorney prior to judge assignment (Section 2.3). We again test this assumption’s validity in Table 2.
To test the validity of the identifying assumptions behind Equations (1) and (2), we measure the linear relationship between case characteristics and a dummy variable for an assigned female judge. Table 2 shows these correlations for each sample of crimes. We find that, after including court district-quarter fixed effects, judge gender is uncorrelated with case characteristics such as whether the defendant was in jail before the case was assigned, whether the defendant was caught in the act by the police and whether the crime is a domestic violence offense. This result suggests that our identifying assumptions are valid.
@@
  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of characteristics of the case. Column (1) shows this relationship for the sample of all physical assault crimes, Column (2) shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years), and Column (3) shows this relationship for the sample of domestic violence crimes. All regressions include quarter-by-court district fixed effects.
@@
We measure the gender conviction rate gap using two empirical strategies. First, we estimate Equation (1) and find that female judges are more likely to convict in domestic violence cases than equally senior men. Second, we estimate Equation (2) and find that the gender conviction rate gap is small or nonexistent for misdemeanors and non-domestic violence physical assaults, whereas the gap becomes large and significant for domestic violence cases.

### v3

@@ Abstract
We investigate the impact of judges’ gender on the outcome of domestic violence cases. Using data from São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p.) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. To show that this decision gap rises due to different gender perspectives about domestic violence and not because female judges are stricter than their male counterparts in all rulings, we compare it against the gender conviction-rate gap in similar types of crime. We find that the gender conviction-rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p. larger) and for other physical assault cases (8 p.p. larger). Furthermore, we find evidence that at least two channels explain this gender conviction-rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria. Lastly, we find that this gender conviction rate has no significant impact on the probability of appeals, ruling reversals or recidivism.
@@ Introduction
Violence against women is frequently perpetrated inside the home by the ones closest to them. The United Nations Office on Drugs and Crime (2023) estimates that 55% of all homicides against women worldwide in 2022 were committed by intimate partners or other family members. As this violence is a criminal offense, key components of the fight against domestic violence are the Criminal Court System and how court judges approach this type of crime (Beecher-Monas, 2001). For instance, the Human Rights Watch (1991) argues that, in the 1980s and early 1990s, Brazilian judges failed to analyze crimes of domestic violence against women in a non-discriminatory manner. [^1] In the years since, Brazilian authorities have made important legal and institutional changes to address domestic violence crimes. The most prominent of these is the Maria da Penha Law, enacted in 2006, which characterizes domestic violence as a specific type of criminal offense. Ferraz and Schiavon (2022) find that this law reduced female homicides due to household aggression by 9 percent, with stronger effects for less educated and black women. Moreover, Kafka et al. (2019) finds that American judges in North Carolina use gender stereotypes to argue that some women want to “game the system” with “frivolous cases” when requesting protective orders against their abusive partners.
@@
Since discriminatory gender stereotypes may vary depending on the decision maker’s individual characteristics, we investigate the impact of judges’ subjectivity on the outcome of domestic violence cases. In particular, we focus on the effect of a judge’s gender on their decisions, documenting a sizeable and significant gender conviction-rate gap specific to domestic violence cases. We also analyze two possible forces driving this difference. At the end, we discuss the consequences of this gender conviction-rate gap beyond the Trial Court, analyzing the outcome of domestic violence cases in the Appeals Court and the recidivism behavior of these cases’ defendants. We find that the Trial Judge’s gender has no effect on these outcomes.
@@
To do this, we collect data from São Paulo, Brazil, between 2010 and 2019 and compare conviction rates by the presiding judge’s gender after controlling for the judge’s career characteristics. Case characteristics are independent of the presiding judge’s gender because cases are randomly allocated to judges within each court district in Brazil. Hence, the gender conviction-rate gap captures different decision criteria by judges of different genders.
@@
In this analysis, we find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to the gender decision gaps for two other types of crime: misdemeanors and physical assault cases. Because we focus on domestic violence crimes that do not result in death or long-term physical impairment, misdemeanors are comparable to domestic violence cases in their sentence severity (according to Brazilian Law). Physical assault cases are comparable to domestic violence cases because both involve some sort of physical aggression.
@@
We find that the gender conviction-rate gap for domestic violence cases is more than twice the size of that of other misdemeanors. Furthermore, while there is a gender conviction-rate gap for domestic violence cases, there are no differences between female and male judges when analyzing other types of physical assault. Consequently, the gender conviction-rate gap for domestic violence cases likely arises due to different perspectives about this type of crime rather than because female judges are, in general, more punitive than their male counterparts.
@@
To investigate if the salience of gender identity matters (representational account), we test if the gender difference in strictness is more pronounced in intimate partner violence cases. To do so, we explore each case’s full sentence and estimate whether female judges are more likely to convict a defendant for a domestic violence offense and simultaneously flag the case as associated with intimate partner violence. We find that the gender conviction-rate gap is 37.5% higher (9 p.p.) in cases in which a female judge may identify more strongly with the victim than a male judge and statistically null in cases that are not flagged as intimate partner violence offenses. This difference is associated with the representational account of the in-group bias because the incidence and tolerance of this type of offense are strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; Tur-Prats, 2019; González and Rodríguez-Planas, 2020).
@@
To test for gender differences in the interpretation of the evidence provided in domestic violence cases (informational account), we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence. We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while they differ when making decisions about cases that rely on more subjective evidence. Hence, female judges process the information contained in domestic violence cases differently from their male peers, as suggested by the informational account of the in-group bias.
Lastly, we discuss the consequences of the gender conviction-rate gap beyond the Trial Court. To do so, we analyze two types of consequences of a Trial Judge’s sentences. First, we look at the outcome of domestic violence cases in the Appeals Court, analyzing whether the Trial Judge’s gender impacts how frequently there is an appeal and how frequently their sentence is reversed. Second, we look at the future behavior of defendants in domestic violence cases, analyzing whether the Trial Judge’s gender impacts the defendants’ criminal recidivism after the final sentence.
The impact of stricter female judges on appeals and ruling reversals in domestic violence cases is theoretically ambiguous. On the one hand, the positive gender conviction-rate gap may lead to female judges facing a larger reversal probability because sentences by stricter judges are more frequently reversed (Possebom, 2023). On the other hand, female judges may exert more effort (informational account) than male judges when analyzing domestic violence cases and, consequently, their rulings might have a smaller probability of being reversed. Importantly, in our empirical context, we find that these two forces balance each other and the Trial Judge’s gender has no impact on the outcome of domestic violence cases in the Appeals Court. For this reason, we conclude that the gender conviction-rate gap does not entail a higher cost in terms of Courts’ resources.
The impact of stricter female judges on recidivism by defendants in domestic violence cases is also theoretically ambiguous. On the one hand, the positive gender conviction-rate gap may lead to female judges facing more recidivism because stricter judges tend to convict the type of defendants whose recidivism increases when punished (Acerenza, Possebom, and Sant’Anna, 2024). On the other hand, female judges may be able to influence defendants’ behaviors in other ways, such as directly signaling that violence against women is not tolerated (representational account). If the typical defendant in a domestic violence case is more likely to recidivate in a similar offense than to commit other types of crime, female judges may lead to lower recidivism. Importantly, in our empirical context, we find that these two forces balance each other and the Trial Judge’s gender has no impact on recidivism. For this reason, we conclude that the gender conviction-rate gap has no negative consequences in terms of future criminal behavior by defendants.
@@
This article contributes to the literature on judicial biases. Harris and Sen (2019) reviews this literature, focusing on ideology, race and gender gaps. Within the studies focusing on in-group bias, many authors find evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Boyd, Epstein, and Martin, 2010; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Similarly to our work, Kruttschnitt and Savolainen (2009), Boyd, Epstein, and Martin (2010), Bielen and Grajzl (2020), and Chen et al. (2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking) in Finland, federal appellate cases in the U.S., civil cases in Belgium, and in cases with female offenders in Kenya, respectively. Our work differs from theirs because we focus on a single common crime type (domestic violence) whose victims are mostly women. Furthermore, while Boyd, Epstein, and Martin (2010) finds that in-group bias in federal appellate cases in the U.S. is consistent with the informational account only, we find that in-group bias in criminal cases in Brazil is consistent with the representational and informational accounts.
More importantly and differently from the previous studies, we not only document in-group bias by female judges in domestic violence cases but also analyze the consequences of this phenomenon beyond the Trial Court. We investigate the effects of the Trial Judge’s gender on the case’s outcomes at the Appeals Court and on the defendant’s future criminal behavior. In particular, we find that the gender conviction-rate gap has no negative consequences in terms of the use of Courts’ resources nor in terms of the defendants’ future criminal behavior.
@@
The paper is organized as follows. Section 2 describes our dataset, while Section 3 explains our empirical strategy. Moreover, Section 4 documents the gender conviction-rate gap, while Section 5 discusses its driving forces. Section 6 analyzes the consequences of the gender conviction-rate gap beyond the Trial Court. Section 7 concludes.
@@
We collect data from all criminal cases brought to the Justice Court System in the State of São Paulo, Brazil, between January 4th, 2010, and December 3rd, 2019. This dataset includes the date each case was assigned to a courtroom, the court district where the crime took place, the judge assigned to the case, the full text of the sentence, a list of the main events of the case, the primary and secondary crime types associated to the case, information on whether the case was analyzed by the Appeals Court, and the full text of the Appeals Court’s ruling.
@@
We use the full-sentence text to construct our main outcome. To do this, we adapt the classification algorithm developed by Possebom (2023) so that it classifies domestic violence case outcomes in a more precise manner. [^3] See Appendix A for an overview of our data-construction process. Differently from Possebom (2023), our main outcome focuses exclusively on conviction decisions. This variable equals one when the judge convicts at least one of the defendants listed in the case of at least one of the charges on the case. As a consequence of conviction, the defendant receives some type of punishment (e.g., incarceration, fine, or community service) and a criminal record.
We also collect information on three types of outcomes that may happen after the Trial Judge’s ruling. First, we measure whether the case was analyzed by the Appeals Court. This event happens if either the defense attorney or the district attorney disagrees with the Trial Judge’s ruling and appeals to a panel of three senior judges. Second, using the classification algorithm proposed by Possebom (2023), we analyze the full text of the Appeals Court’s ruling to measure whether the Trial Judge’s ruling was reversed or not. [^4] See Appendix A for an overview of our data-construction process. Third, we observe whether the defendant recidivated after their final sentence’s date and the type of the new criminal offense. To measure recidivism, we check whether the defendant’s name appears in any criminal case within two years after the final sentence date. [^5] See Appendix A for an overview of our data-construction process.
Furthermore, the district attorney’s office defines each case’s primary and secondary crime type prior to judge assignment. We construct three samples based on this classification.
@@
Second, our ‘physical assault’ sample contains cases whose primary type is physical assault. These cases are associated with other forms of assault, such as street fights, bar fights, or football hooliganism. We call ‘other physical assault’ the physical assault cases that do not involve domestic violence.
Third, our ‘misdemeanors’ sample contains cases that carry a maximum incarceration sentence of less than four years. This threshold is relevant because, according to Brazilian Law, being convicted of these crime types leads to a criminal record but not to incarceration: sentences must be converted into fines or community service. As discussed by Possebom (2023) and Acerenza, Possebom, and Sant’Anna (2024), most misdemeanor crimes in Brazil are theft offenses. We call ‘other misdemeanors’ the misdemeanor offenses unrelated to domestic violence.
@@
The productivity data from the São Paulo Court of Justice website contains records on the total amount of work done by each judge in any given month and court district from January 2011 to December 2019.777For example, it measures the number of hearings or disposed cases per month for each judge in each court district. We use this information to keep track of which judges are active in each court district-month pair. [^8] Months are the natural unit of time for keeping track of activity and work volume because salaries in Brazil are paid monthly. Furthermore, most hiring, reallocation, and termination takes effect at the beginning of each month.
@@
Our final dataset is at the criminal-case level. It is crucial for our identification strategy that the gender of the judge is randomly assigned and correctly identified for each case. To enforce these two characteristics, we adopt four steps.
@@
Third, since we compare decisions made by female and male judges, we require the judge’s gender to be randomly allocated to cases.101010 We measure judges’ gender using a two-step procedure. First, we apply the classification algorithm proposed by the R package genderBR, which uses the share of men and women with each name in the 2010 Brazilian Census. Since a few names are not commonly associated with a specific gender, we visit the São Paulo Court System’s website, check the photos of the remaining judges and manually classify their gender. The same classification procedure was used by Laneuville (2024). In other words, the ex-ante probability of a female judge being assigned to each case must be different from zero or one, which would not be the case in courts without judges of both genders. To enforce this restriction, we limit the sample to cases from court district-month pairs with at least one male and one female active judge in the productivity dataset.
Last, we need to ensure that the assigned judge’s gender is determined correctly. To do this, we focus on cases allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment. [^11] Cases are randomly allocated to courtrooms instead of directly to individual judges. If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
Using this final sample, we compute summary statistics for cases’ outcomes and crime types. Table 1 displays cases’ outcomes in the columns and crime types in the rows. It also reports the share of female judges analyzing each type of crime. Domestic violence cases are more frequently convicted than other physical assault cases but are less frequently convicted than other misdemeanor cases. We also note that the share of female judges is similar across crime types and that there are 270 female judges and 425 male judges in our entire sample.
Table 1: Cases’ Outcomes and Share of Female Judges by Crime Type in %
Crime Conviction Female Obs
@@
DV 32.5 32.1 4,358
Other Mis 36.9 36.4 39,815
Other PA 27.2 38.1 1,648
  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Other Mis), and physical assault cases that are not domestic violence (Other PA). The first column indicates the frequency of cases whose final ruling is conviction. The conviction dummy equals one when at least one defendant is convicted in at least one charge. The second column shows the share of female judges in each sample, while the last column shows the size of each sample.
@@
Empirical Strategy
We measure the gender conviction-rate gap for domestic violence cases by comparing the conviction probability for similar cases allocated to female and male judges. We do this by running a linear regression model on the sample of domestic violence crimes:
@@
where $\text{Convicted}_{it}$ indicates whether the cases resulted in a conviction, $\text{Female Judge}_{it}$ indicates whether the case was assigned to a female judge, $X_{it}$ are covariates including pre-determined case characteristics (an indicator for whether the defendant was already in jail and an indicator for whether the police caught them committing the crime) and pre-determined judge characteristics (career stage, total experience, and experience in their current career stage as of the day the case was assigned), and “$\text{Court District}_{i}\times\text{Quarter}_{t}$” is a full set of court district-quarter pair fixed effects.
@@
Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence. To measure whether this conviction-rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction-rate gap against this gap for similar crime types. Now, our sample consists of domestic violence cases and one type of comparable criminal case. We estimate the following linear regression model:
@@
where “$\text{Domestic Violence}_{it}$” indicates whether the case is associated with domestic violence. We use two categories of comparable crime types: (i) offenses with similar sentences (misdemeanors) and (ii) offenses of a similar nature (physical assault).
In Equation (2), our parameter of interest is $\beta_{3}$. This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of crimes of comparable severity (coefficient $\beta_{2}$) and for the possibility that female judges are stricter than male judges in general (coefficient $\beta_{1}$). Consequently, the $\beta_{3}$ coefficient captures the part of the gender conviction-rate gap that is likely due to different gender perspectives about domestic violence cases specifically. For this reason, we interpret this coefficient as capturing a type of in-group bias according to the definition given by Shayo and Zussman (2011). We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) may act differently from members of the other group (male judges) when analyzing this type of crime in particular.
This interpretation for the $\beta_{3}$ coefficient is valid if the judge’s gender is independent of the case’s potential outcome after conditioning on the type of crime. This exogeneity assumption is again valid according to Brazilian Law since cases are randomly allocated to judges within court districts, and crime types are determined by the district attorney prior to judge assignment (Section 2.3). We again test this assumption’s validity in Table 2.
Validity Tests
To test the validity of the identifying assumptions behind Equations (1) and (2), we measure the linear relationship between case characteristics and a dummy variable for an assigned female judge. Table 2 shows these correlations for each sample of crimes. We find that, after including court district-quarter fixed effects, judge gender is uncorrelated with case characteristics such as whether the defendant was in jail before the case was assigned, whether the defendant was caught in the act by the police and whether the crime is a domestic violence offense. These results suggest that our identifying assumptions are valid.
@@
  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of case characteristics. Column (1) shows this relationship for the sample of all physical assault crimes, Column (2) shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years), and Column (3) shows this relationship for the sample of domestic violence crimes. All regressions include quarter-by-court district fixed effects.
@@
We measure the gender conviction-rate gap using two empirical strategies. First, we estimate Equation (1) and find that female judges are more likely to convict in domestic violence cases than equally senior men. Second, we estimate Equation (2) and find that the gender conviction-rate gap is small or nonexistent for misdemeanors and non-domestic violence physical assaults, whereas the gap becomes large and significant for domestic violence cases.